Arlington and Vernonville are 144 miles apart. A car leaves Arlington traveling towards veronville, and another car leaves veronville at the same time, traveling towards Arlington. The car leaving Arlington averages 10 miles per hour more than the other, and they meet after 1 hour and 36 minutes. What are the average speeds of the drivers?
The car leaving Arlington averages mph?
The car leaving Veronville averages mph?

Answer by mananth(16946)

Arlington and Vernonville are 144 miles apart. A car leaves Arlington traveling towards veronville, and another car leaves veronville at the same time, traveling towards Arlington. The car leaving Arlington averages 10 miles per hour more than the other, and they meet after 1 hour and 36 minutes. What are the average speeds of the drivers?

Arlington bound car speed = x mph
Vernonville bound car = x+10 mph
distance = 144 miles
...
since they are travelling towards each other speed = 2x+10 mph
time = distance/speed
3/5 = 144/(2x+10)
3*(2x+10)=144*5
6x+30 = 720
6x=720-30
6x=690
x= 115 mph Arlington bound car speed
Vernonville bound car speed = 125 mph
...
m.ananth@hotmail.ca
The car leaving Arlington averages mph?
125 mph
The car leaving Veronville averages mph?
115mph